OpenStudy (anonymous):

The radioactive substance strontium -90 has a half-life of 28 years. In other words, it takes 28 years for half of a given quantity of strontium -90 to decay to a non radioactive substance. The amount of radioactive strontium -90 still present after t years is modeled by the expression 80*2^-t/28 grams. Evaluate the expression for t=0, t=28, and t=56. What does each value of the expression represent?

OpenStudy (anonymous):

First of all the half life is the amount of time it takes for a substance to decay to half of its original amount. \[80\times2^{-t/28}\] All you need to do is plug in t=0, t=28 and t=56. Then what does it mean when we evaluated at time zero? how about after 28 years? after 56?

OpenStudy (anonymous):

how do you plug it in cause iv never did this b4

OpenStudy (anonymous):

Solving for zero we just replace t with zero which gives\[80\times2^{-\frac{ -(0) }{ 28 }}\] which simplifies to \[80\times2^{0}\] and number to the power 0 is 1. So we have \[80\times1=80\] What does this mean? After no time has passed we have 80 units of this strontium 90

OpenStudy (anonymous):

do you have a calculator that can calculate exponents?

OpenStudy (anonymous):

I have no clue this is my first time seeing a question like this

OpenStudy (anonymous):

What kind of calculator do you have?

OpenStudy (anonymous):

i think its a scientific one

OpenStudy (anonymous):

ok good. Do you see the button that looks like this\[x ^{y} \] or ^

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k thats the button that does exponents. so to evaluate at t=28 your going to plug into your calculator kind of like this: 80 X 2^(-28/28)

OpenStudy (anonymous):

you should get 40

OpenStudy (anonymous):

cause 2^(-28/28) is 2^(-1) which is 1/2 and 80 X 1/2= 40

OpenStudy (anonymous):

yep thats what i got

OpenStudy (anonymous):

good now try for t=56 and see what you get

OpenStudy (anonymous):

so instead of 28 put 56?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

its 40 also?

OpenStudy (anonymous):

make sure you only plug in 56 for t. so 80X2^(-t/28) will be 80X2^(-56/28)

OpenStudy (anonymous):

oh ok i put 56 twice instead of 28 hold on

OpenStudy (anonymous):

its 20

OpenStudy (anonymous):

yup now you just need to think about what each value represents

OpenStudy (anonymous):

ok umm how do you do that lol sorry im a little slow :)

OpenStudy (anonymous):

so at t=0 that means no time has passed and we have 80 units of the substance so 80 is our original amount. at t=28 one half life has passed because the half life of theis substance in 28 years so at t=28 half of the original amount is present. 40 right? then at t=56 two half lives have passed because 28+28=56 and so 1/2*1/2=1/4 so we should have 1/4 of the original amount and we saw that at t=56 there were 20 units of the original amount. and 80X(1/4)=20.

OpenStudy (anonymous):

so the value that they represent is 20?

OpenStudy (anonymous):

each time value (t=0, t=28 and t=56) represents something different. t=0 is the original amount t=28 is half of the original amount and t=56 is half of half of the original amount

OpenStudy (anonymous):

oh ok thank you so much @Mr.Me